OpenStudy (anonymous):

OKAY this one is really confusing to me... Subtract, simplify by removing a factor of 1 when possible 11fc/f^2-c^2 - f-c/f+c

OpenStudy (anonymous):

always start by finding a common denominator.

OpenStudy (anonymous):

11 and 2 are not common.. so 1??

OpenStudy (anonymous):

any factorization will be done after you balance the denominators. So, you might not see it immediately. Math problems will often look crazy until you make some adjustments.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{11fc}{f^2-c^2} - \frac{f-c}{f+c}\] \[\Large \frac{11fc}{(f-c)(f+c)} - \frac{f-c}{f+c}\] \[\Large \frac{11fc}{(f-c)(f+c)} - \frac{(f-c)(f-c)}{(f-c)(f+c)}\] \[\Large \frac{11fc-(f-c)(f-c)}{(f-c)(f+c)}\] \[\Large \frac{11fc-(f^2-2fc+c^2)}{(f-c)(f+c)}\] \[\Large \frac{11fc-f^2+2fc-c^2}{(f-c)(f+c)}\] \[\Large \frac{-f^2+13fc-c^2}{(f-c)(f+c)}\] \[\Large \frac{-(f^2-13fc+c^2)}{(f-c)(f+c)}\] \[\Large -\frac{f^2-13fc+c^2}{(f-c)(f+c)}\] \[\Large -\frac{f^2-13fc+c^2}{f^2-c^2}\] So \[\Large \frac{11fc}{f^2-c^2} - \frac{f-c}{f+c}=-\frac{f^2-13fc+c^2}{f^2-c^2}\]

OpenStudy (anonymous):

there you go, that dude is CRAZY!

jimthompson5910 (jim_thompson5910):

The key here is to get all denominators equal to (f-c)(f+c). From there, you can combine the fractions and simplify.

OpenStudy (anonymous):

yeah Jim is a beast I'm in love with him <3 lol

OpenStudy (anonymous):

it says its not rifght :(

jimthompson5910 (jim_thompson5910):

then try \[\Large \frac{-f^2+13fc-c^2}{f^2-c^2}\] if that doesn't work, then try \[\Large \frac{-f^2+13fc-c^2}{(f-c)(f+c)}\] the computer is very picky and will only accept one answer even though there are multiple ways of saying the same thing

OpenStudy (anonymous):

nvm i typed it in wrong!!

jimthompson5910 (jim_thompson5910):

ah gotcha

OpenStudy (anonymous):

well nvm its still not wright lol

OpenStudy (anonymous):

it's okay though :D

jimthompson5910 (jim_thompson5910):

well it's one of forms listed above

OpenStudy (anonymous):

i didnt put parenthesis.. thats why

jimthompson5910 (jim_thompson5910):

often, even the order matters (even though it shouldn't)

jimthompson5910 (jim_thompson5910):

i see